OpenStudy (anonymous):

The initial population is 91. The population doubles every 14 years. What is the equation for the population after t number of years?

OpenStudy (saifoo.khan):

Geometric progression

OpenStudy (saifoo.khan):

Use for formula, \[\Large T_n = ar^{n-1}\]

OpenStudy (saifoo.khan):

Where a is initial popl. i.e. 91.

OpenStudy (anonymous):

every 14 years means use \(A_0\times 2^{\frac{t}{14}}\)

OpenStudy (anonymous):

n(t)=n0e^kt

OpenStudy (anonymous):

n0=initial population

OpenStudy (anonymous):

find k then use k to find t

hero (hero):

There's probably more that one approach to this, but I think @cammyabbo has the more useful one.

OpenStudy (anonymous):

91(2)^(t/14)?

OpenStudy (anonymous):

once you have e, you take the ln of both sides to bring down your exponent

OpenStudy (anonymous):

182=91e^k(2)

OpenStudy (anonymous):

becomes 91=e^k(2)

OpenStudy (anonymous):

to hell with the e, too much work

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

you are told the doubling time, use 2 as the base

OpenStudy (anonymous):

2=e^k(2)

OpenStudy (anonymous):

take the ln of both sides

OpenStudy (anonymous):

ln 2=k(2)

OpenStudy (anonymous):

k=ln 2/2 plug back into original equation to find t

hero (hero):

I don't remember you quoting a general formula @satellite73

OpenStudy (anonymous):

You all have royally confused me. lol

OpenStudy (anonymous):

??

OpenStudy (anonymous):

follow my posts

OpenStudy (anonymous):

you are told the doubling time. doubles every 14 hours right? so it look like this at time 0 you have 91, at time 14 you have \(91\times 2\) at time 28 you have \(91\times 4=91\times 2^2\) and in general you have \(91\times 2^x\)

OpenStudy (anonymous):

now since it doubles every 14 year you know that the exponent should look like \(\frac{t}{14}\) so that if t is a multiple of 14 you get just what you want. now you can make t any value you like so general formula is \[91\times 2^{\frac{t}{14}}\]

hero (hero):

That's not general enough. I'm referring to a global general math formula that you can apply to any problem. I know you're the man on here, but there's a reason why general formulas were created.

OpenStudy (anonymous):

there is no reason not to use 2 as a base if you are doubling

hero (hero):

You always seem to neglect posting such general formulas.

OpenStudy (anonymous):

ok suppose i have a population that increases by 20% every 3 hours. and lets say i start with P then instantly i know i can model this as \[P(1.2)^{\frac{t}{3}}\] what more do i need?

OpenStudy (anonymous):

if if the half life is h, i use \[A_0(\frac{1}{2})^{\frac{t}{h}}\] done

OpenStudy (anonymous):

this going back and forth with e is math teacher nonsense inflicted on students for no good reason at least not at this level

hero (hero):

You could have posted that general formula at the beginning of your explanation. It helps other users understand your basis of approach

OpenStudy (anonymous):

is \[A_0(\frac{1}{2})^{\frac{t}{h}}\] general ?

OpenStudy (anonymous):

but this calls for going straight to the answer using your eyes and head, not a formula doubling time is 14 years, initial population is 91, use \[91\times 2^{\frac{t}{14}}\]an explanation might be needed but a general formula is not

hero (hero):

I don't know if it is or not, but it's better than just presenting some equation where it looks like you just randomly posted some numbers from the problem and threw an equation together. As far as using eyes and head, you're losing sight on the fact that not every user is as familiar with such things as you are.

hero (hero):

You should always post with other users in mind.

hero (hero):

That which is so obvious to you may be so confusing for someone else.

OpenStudy (anonymous):

@Aprilmbug what numbers were you given? a) doubles use the number 2 as a base b) time: 14 years c) initial population, 91 formula : \(91\times 2^{\frac{t}{14}}\) what could be more straightforward?

OpenStudy (anonymous):

randomly posted numbers? i respectfully disagreed. i used only the numbers written above

hero (hero):

I could write 2 x 14 divided by 91 = t but that would be misleading, wouldn't it?

OpenStudy (anonymous):

compare my formula with the one with the log and the e and all that (which is of course correct, although there is a minor mistake, it should be \(2=e^{14k}\) but no matter

OpenStudy (anonymous):

well the math teachers teach us in these ways

OpenStudy (anonymous):

im not a math wiz, i have to follow general rules and formulas, otherwise i mess up

OpenStudy (anonymous):

and what do you see? a bunch of steps to change the base from 2, which it is because you are givne the doubling time, to the base of e using logs for god sake. why?

OpenStudy (anonymous):

that is one way of solving the problem, i hope it helped

OpenStudy (anonymous):

Yeahh I just went to wolfram alpha. but thanks guys